Balance chemical reaction
MgO + HCL

Cr(OH)3

Cu(NO3)2 + Mg
1
Expert's answer
2017-10-02T11:42:07-0400
MgO + 2HCL =MgCl2 + H2O

2Cr(OH)3 (t)= Cr2O3 + 3H2O

Cu(NO3)2 + Mg = Mg(NO3)2 + Cu